OpenStudy (anonymous):

What is y=-2/3x+4 in standard form?

OpenStudy (anonymous):

helpppppppp

OpenStudy (anonymous):

Given equation is \(\large y=\frac{-2}{3}x+4\) first we bring it to the standard form by taking x to the left of equal sign,this becomes now as, \(\large \frac{2}{3}x+y=4 \) now the coefficient of x is not an integer, we need to make it an integer, for that multiply the whole equation by 3, then the equation becomes.. \(\large 2x+3y=12\) this equation is now in standard form.